OpenStudy (anonymous):

Help please: ß=11pi/2 It says to find the value of the six trig functions My teacher also has that it's equal to 3pi/2 which I don't understand

OpenStudy (diamondboy):

the question is not appearing properly

OpenStudy (anonymous):

\[\beta=11/2\pi\] It says to find the exact value for each of the six trig functions @diamondboy

OpenStudy (diamondboy):

how far have u gone?

OpenStudy (anonymous):

I wrote the formula down that's it

OpenStudy (diamondboy):

did ur teacher say \[11/2\pi = 3\pi/2\]

OpenStudy (anonymous):

Yes

OpenStudy (diamondboy):

ok

OpenStudy (diamondboy):

lets walk through this together, the 6 trig functions are sin,cosine,tangent,cotangent,cosecant and secant

OpenStudy (diamondboy):

do u knw ur unit circle?

OpenStudy (anonymous):

I don't know it well

OpenStudy (diamondboy):